what is the aspect ratio of television screen is the ratio of the measure of horizontal length to the measure of the vertical length. Find the horizontal length and vertical height of a television screen with an aspect ratio of 4 to 5 and a diagonal of 50 inches.

Answer by stanbon(75887)

what is the aspect ratio of television screen is the ratio of the measure of horizontal length to the measure of the vertical length. Find the horizontal length and vertical height of a television screen with an aspect ratio of 4 to 5 and a diagonal of 50 inches.
horizontal length = 4x
vertical length = 5x
----
Equation:
(4x)^2 + (5x)^2 = 50^2
16x^2 + 25x^2 = 50^2
41x^2 = 2500
x = 50/sqrt(41)
horizontal = 200/sqrt(41)"
vertical = 250/sqrt(41)"
-------------
Cheers,
Stan H.
------------